Determine if the lines 9x - 2y = 16 and 2x - 9y = -18 are parallel, perpendicular, or neither
Answer by RAY100(1637)

one way to check slopes is to have equations in slope intercept form ( y=mx +b )
.
(1) 9x -2y =16
-2y =-9x +16
y =( 9/2) x -8
.
(2) 2x -9y = -18
-9y = -2x -18
y = (2/9) x +2
.
If parallel, m1 =m2,,,,no
.
if perpendicular m1= - 1/m2,,,,,no
.
therefore other,,in fact intersect at ( 2.34, 2.52)